OpenStudy (anonymous):

what is 0.26 in a fraction

sammixboo (sammixboo):

Well, \(\rm \color{#20bd23}{0.26}\) is equal to 26%, and 26% means 26 per 100, so we would have \(\rm \color{#20bd23}{\dfrac{26}{100}}\), do you know how to simplify fractions?

OpenStudy (anonymous):

26/ 100=13/50

OpenStudy (ageta):

Step 1: 0.26 = 26⁄100 Step 2: Simplify 26⁄100 = 13⁄50

OpenStudy (ageta):

13/50

OpenStudy (ageta):

does that help

OpenStudy (ageta):

tnxs

OpenStudy (anonymous):

13/50

OpenStudy (anonymous):

yes